### v1

# Lower bounds for Seshadri constants on blow ups of $\mathbb{P}^{2}$

Abstract.
Let $\pi:X_{r}\rightarrow\mathbb{P}^{2}$ be a blow up of $\mathbb{P}^{2}$ at $r$ distinct points $p_{1},p_{2},\dots,p_{r}$. We study lower bounds for Seshadri constants of ample line bundles on $X_{r}$. First, we consider the case when $p_{1},p_{2},\dots,p_{r}$ are on a curve of degree $d\leq 3$. Then we assume that the points are very general and show that $\varepsilon(X_{r},L,x)\geq 1$ for any ample line bundle $L$ and any $x\in X_{r}$ if the Strong SHGH conjecture is true. We explore the relation between bounded negativity and Seshadri constants and study the Seshadri function on $X_{r}$.
Key words and phrases:
Seshadri constant, Bounded Negativity, Seshadri function
2020 Mathematics Subject Classification:
14C20, 14E05, 14H50
Author was partially supported by a grant from Infosys Foundation
1. Introduction
Seshadri constants are used to study the local positivity of ample line bundles on a projective variety. They were introduced by Demailly in [3]. Consider a smooth complex projective variety $X$. The Seshadri constant of a nef line bundle $L$ on $X$ at a point $x\in X$ is defined as follows:
$\varepsilon(X,L,x)=\inf\bigg{\{}\frac{L\cdot C}{\text{mult}_{x}C}~{}~{}\bigg{|}~{}~{}C\text{ is a reduced irreducible curve in $X$ passing through $x$}\bigg{\}}.$
There is an alternate definition for the same, given as follows:
Consider the blow up $\pi_{x}:\widetilde{X}(x)\rightarrow X$ of $X$ at $x$. Let $E=\pi_{x}^{-1}(\{x\})$ denote the exceptional divisor. Then
$\varepsilon(X,L,x)=\sup\{t\in\mathbb{R}~{}~{}|~{}~{}\pi_{x}^{\ast}L-tE\text{ is a nef bundle on }\widetilde{X}(x)\}.$
This definition enables us to conclude that $\varepsilon(X,L,x)\leq\sqrt{L^{2}}$. This inequality motivates us to give following definition:
Definition 1.1.
Let $L$ be an ample line budle on a smooth projective surface $X$ and let $x\in X$. We say that $L$ is submaximal at $x$ if $\varepsilon(X,L,x)<\sqrt{L^{2}}$.
If $L$ is submaximal at $x$ then there exists a reduced and irreducible curve $C$ passing through $x$ such that $\varepsilon(X,L,x)=\frac{L\cdot C}{\text{mult}_{x}C}$. Such curves are called Seshadri curves of $L$ at $x$. If $\widetilde{C}$ is the strict transform of a Seshadri curve $C$ on $\widetilde{X}(x)$, then by the Hodge index theorem we have $\widetilde{C}^{2}<0$. A reduced irreducible curve $C$ on $X$ is said to be a $(-m)$ curve if $C^{2}=-m$ and $K_{X}\cdot C=m-2$, where $K_{X}$ is the canonical line bundle of $X$.
Further, we will define two more constants on a surface $X$ which are related to the Seshadri constant of an ample line bundle at a point. Firstly, we will define Seshadri constant of $X$ at point $x\in X$, denoted by $\varepsilon(X,x)$, as follows:
$\varepsilon(X,x)=\inf_{L\text{ ample }}\varepsilon(X,L,x).$
Secondly, the Seshadri constant of a surface $X$ is defined to be:
$\varepsilon(X)=\inf_{x\in X}\varepsilon(X,x).$
More details about Seshadri constants and local positivity can be found in [16, Section 5] and [1]. One of the main problems in the area of Seshadri constants is to give lower bounds for $\varepsilon(X)$. Ein and Lazarsfeld proved that $\varepsilon(X,x)\geq 1$ for very general points $x\in X$; see [5, Theorem]. But Miranda showed that for every $\delta>0$ there is a surface $X$ and an ample line bundle $L$ on $X$ such that $\varepsilon(X,L,x)\leq\delta$ at a point $x\in X$ (see [5, Example 3.1] and [16, Example 5.2.1]). In this article, we will address this question when $X$ is a blow up of $\mathbb{P}^{2}$ at finitely many points.
Let $\pi:X_{r}\rightarrow\mathbb{P}^{2}$ be the blow up of $\mathbb{P}^{2}$ at $r$ distinct points $p_{1},p_{2},\dots,p_{r}$. Define $H$ to be the pullback of a line not passing through any $p_{i}$ and let $E_{i}=\pi^{-1}(\{p_{i}\})$ denote the exceptional divisor corresponding to $p_{i}$, for all $1\leq i\leq r$ Then the linear equivalence classes of $H,E_{1},E_{2},\dots,E_{r}$ form a basis for the class group $\text{Cl}(X_{r})$ of $X_{r}$. We have
$H^{2}=-E_{i}^{2}=1\text{ and }H\cdot E_{i}=E_{i}\cdot E_{j}=0\text{ for all }i\neq j.$
With this basis, the canonical bundle of $X_{r}$ is $K_{X_{r}}:=-3H+E_{1}+E_{2}+\dots+E_{r}$. As the Seshadri constant at point $x$ in $X_{r}$ has an alternative definition as defined above, we will often consider the blow up $\pi_{x}:\widetilde{X}_{r}(x)\to X_{r}$ at $x$. By abuse of notation, $E_{i}$ together with $E_{x}:=\pi_{x}^{-1}(\{x\})$ will denote the exceptional divisors for $\widetilde{X}_{r}(x)$.
For a smooth projective surface $X$, in a recent work, Schmidt [18] defined a function called the Seshadri function on $\text{Nef}(X)$,
$\varepsilon_{x}:\text{Nef}(X)\rightarrow\mathbb{R}$
defined as follows:
$\varepsilon_{x}(L)=\varepsilon(X,L,x).$
He proved that this function is continuous and is piecewise linear in a neighborhood of an ample line $L$ in $\text{Nef}(X_{r})$ with $\varepsilon(X,L,x)<\sqrt{L^{2}}$. He also showed that
$W_{x}=\{L\in\text{Nef}(X)~{}~{}|~{}~{}\varepsilon(X,L,x)<\sqrt{L^{2}}\}$
is nowhere dense when $X$ is an abelian surface [18, Lemma 4.1]. In last section we study the Seshadri function on $X_{r}$.
To study the Seshadri constant we will look at linear systems on $\mathbb{P}^{2}$ as well as on $X_{r}$. There are many conjectures proposed by several authors in this direction. We briefly recall these below.
Consider a linear system $\mathcal{L}=|dH-m_{1}E_{1}-m_{2}E_{2}-\dots-m_{r}E_{r}|$ for some non-negative integers $m_{1},m_{2},\dots,m_{r}$, and $d\geq 1$ on a blow up $X_{r}$ of $\mathbb{P}^{2}$ at very general points $p_{1},\dots,p_{r}$. We say that $\mathcal{L}$ is non-special if dim $\mathcal{L}$ is equal to the expected dimension of $\mathcal{L}$, denoted by edim $\mathcal{L}$ and defined as
edim =$\text{edim }\mathcal{L}=\max\Bigg{\{}{d+2\choose 2}-\displaystyle\sum_{i=1}^{r}{m_{i}+1\choose 2}-1,-1\Bigg{\}}.$
Following is one of the main conjectures in this area, independently proposed by Segre[19], Harbourne[12], Gimigliano[8] and Hirschowitz[15].
Conjecture 1.2 (SHGH Conjecture).
If $\mathcal{L}$ is special, then every divisor in $\mathcal{L}$ is non-reduced.
The above version was given by Segre in 1964. There are several versions given by other authors also. It is known that SHGH is true when $r\leq 8$. We also recall the following conjecture which is stronger than the SHGH conjecture, known as Strong SHGH conjecture or Generalized SHGH conjecture [6, Conjecture 3.6]. This was proposed by Ł. Farnik, K. Hanumanthu, J. Huizenga D. Schmitz and T. Szemberg in 2020 in order to exhibit the existence of ample line bundles with an irrational Seshadri constant at every point of a general blow up of $\mathbb{P}^{2}$.
Conjecture 1.3 (Generalized SHGH or Strong SHGH).
Let $X_{r}$ be a blow up of $r$ very general points of $\mathbb{P}^{2}$ and let $d\geq 1$, $t\geq 1$ and $m_{1},m_{2},\dots,m_{r}\geq 0$ be integers such that
(${d+2\choose 2}-\displaystyle\sum_{i=1}^{r}{m_{i}+1\choose 2}\leq\max\Bigg{\{}{t+1\choose 2}-2,0\Bigg{\}}.$
Then any curve $C\in|dH-m_{1}E_{1}-\dots-m_{r}E_{r}|$ which has a point of multiplicity t is non-reduced.
In particular, the case $t=1$ is equivalent to the SHGH conjecture. The following conjecture follows from the SHGH conjecture.
Conjecture 1.4 (Weak SHGH).
Let $X_{r}$ be a blow up of $r$ very general points of $\mathbb{P}^{2}$. If $C$ is an irreducible and reduced curve in $X_{r}$ such that $C^{2}<0$, then $C$ is a $(-1)$-curve.
For $r\geq 9$, Conjecture 1.2 and Conjecture 1.4 are open and Conjecture 1.3 is open even for $r\geq 1$.
We say that a surface $X$ has bounded negativity if there exists an integer $b(X)$ such that $C^{2}\geq-b(X)$ for every irreducible and reduced curve $C$ on $X$. Over a positive characteristic field, there exists a surface without bounded negativity [14, Exercise V.1.10]. The following conjecture is open for surfaces over a field of characteristic zero.
Conjecture 1.5 (Bounded Negativity Conjecture).
For each smooth complex projective surface $X$ there exists an integer $b(X)>0$ such that
$C^{2}\geq-b(X)$
for every irreducible reduced curve $C\subset X$.
As mentioned above we are interested in lower bounds of $\varepsilon(X_{r})$. We give lower bounds when the blown up points are on a curve of degree $d\leq 3$ (see Theorem 2.1), when $r\leq 8$ (see Theorem 3.3 and Theorem 3.6). For arbitrary $r$, when the blown up points are very general, we show that $\varepsilon(X_{r})\geq 1$ assuming Conjecture 1.4 is true (see Theorem 4.2).
In Section 5, we study the relation between bounded negativity and Seshadri constants. We give a lower bound for Seshadri constants assuming Conjecture 1.5. In Section 6, we study the Seshadri function of $X_{r}$. Further, we discuss the submaximality of ample line bundles on $X_{r}$.
2. Blow ups of points on a curve
Let $X_{r}$ denote a blow up of $\mathbb{P}^{2}$ at $r$ points on a curve of degree $d$ in $\mathbb{P}^{2}$. In this section, we will exhibit a lower bound for the Seshadri constant of ample line bundles at any point of $X_{r}$.
Theorem 2.1.
Let $p_{1},p_{2},\dots,p_{r}$ be r distinct points lying on a curve $C\subseteq\mathbb{P}^{2}$ of degree $d\leq 3$. Let $X_{r}$ the denote the blow up of $\mathbb{P}^{2}$ at $p_{1},p_{2},\dots,p_{r}$. Then for any ample line bundle $L$ on $X_{r}$ and $x\in X_{r}$, we have
$\varepsilon(X_{r},L,x)\geq\begin{cases}1\text{ if $d=1,2$ or $d=3$ and $L\cdot(-K_{X_{r}})\geq 2$}\\ \frac{1}{2}\text{ if $d=3$ and $L\cdot(-K_{X_{r}})=1$}\end{cases}$
Proof.
By [12, Theorem III (a)], an ample line bundle on a smooth anticanonical rational surface $X$ is base point free if $L\cdot(-K_{X})\geq 2$. It is well-known that for an ample base point free line bundle $L$ on a surface $X$, $\varepsilon(X,L,x)\geq 1$ for every $x\in X$ [16, Example 5.1.18]. First, we consider the case $d=3$. If $L\cdot(-K_{X_{r}})\geq 2$ then $\varepsilon(X_{r},L,x)\geq 1$. If $L\cdot(-K_{X_{r}})=1$, $2L$ will be an ample base point free line bundle, which gives $\varepsilon(X_{r},2L,x)\geq 1$. This is equivalent to $\varepsilon(X_{r},L,x)\geq\frac{1}{2}$.
To finish, assume that $d=1$ or $d=2$. Let $L$ be an ample line bundle on $X_{r}$ and let $\widetilde{C}$ be the strict transform of $C$ on $X_{r}$. Since $L$ is ample we have $L\cdot\widetilde{C}\geq 1$ and $L\cdot H\geq 1$. As $-K_{X_{r}}=\widetilde{C}+(3-d)H$ and $d\leq 2$, it follows that $L\cdot(-K_{X_{r}})\geq 2$. ∎
Now we consider the case of points lying on a curve of arbitrary degree.
Proposition 2.2.
Let $C$ be an irreducible and reduced curve of degree $d$ in $\mathbb{P}^{2}$ and $X_{r}$ the blow up of $r$ distinct smooth points $p_{1},p_{2},\dots,p_{r}\in C$. If $L=eH-n(E_{1}+\dots E_{r})$ is an ample line bundle on $X_{r}$ then,
$\varepsilon(X_{r},L)\geq\frac{1}{r}~{}\;~{}~{}\;\forall r\geq d^{2}+1.$
Proof.
Let  $\widetilde{C}=dH-E_{1}-\dots-E_{r}$ be the strict transform of $C$. As above, it is enough to show that $rL$ is globally generated. By [9, Theorem 3.2] for $r\geq d^{2}+1$, $rL$ is globally generated if $(re+3)d>r(rm+1)$. This is equivalent to $r(ed-rm)>r-3d$. This is true since $L\cdot\widetilde{C}=ed-rm\geq 1$. ∎
3. Seshadri constants on Blow ups of $r\leq 8$ distinct points
In this section, we consider blow ups of $\mathbb{P}^{2}$ at $r\leq 8$ points and give lower bounds for Seshadri constants. These partially follow from Theorem 2.1. Throughout this section, we will assume that $r\leq 8$ and $X_{r}$ denotes the blow up of $\mathbb{P}^{2}$ at $r$ distinct points. We recall the following proposition.
Proposition 3.1.
[11, Proposition 2.5] Let $X_{r}$ be the blow up of $r$ distinct points in $\mathbb{P}^{2}$ and let $L$ denote an ample line bundle on $X_{r}$. Assume $r\leq 8$, and for $x\in X_{r}$, let $\widetilde{X}_{r}(x)$ denote the blow up of $X_{r}$ at $x$. Then
$\varepsilon(X_{r},L,x)=\inf\bigg{\{}\frac{L\cdot C}{\rm{mult}_{x}C}~{}~{}\bigg{|}~{}~{}\widetilde{C}\in\Gamma\bigg{\}},$
where $\Gamma$ is the set of all $(-1)$-curves, $(-2)$-curves, fixed components of $\big{|}-K_{\widetilde{X}_{r}(x)}\big{|}$ or reduced irreducible curves in the linear system $\big{|}-K_{\widetilde{X}_{r}(x)}\big{|}$ on $\widetilde{X}_{r}(x)$.
Remark 3.2.
If $C$ is a Seshadri curve in $X_{r}$ for a line bundle at a point $x\in X_{r}$, then $\widetilde{C}^{2}<0$, where $\widetilde{C}$ is the strict transform of $C$ on $\widetilde{X}_{r}(x)$. Note that by [17, Proposition 4.1 (ii)], for $r\leq 8$, $\widetilde{C}$ is either $(-1)$-curve, $(-2)$-curve, or a fixed component of $\big{|}-K_{\widetilde{X}_{r}(x)}\big{|}$.
Proposition 3.3.
Let $X$ be a smooth anticanonical surface (i.e., $-K_{X}$ is an effective divisor). If $(-K_{X})^{2}\geq 2$ then for every $x\in X$,
$\varepsilon(X,L,x)\geq 1$
for any ample line bundle $L$ on $X$.
Proof.
As an application of the Hodge index theorem we have,
$(-K_{X}\cdot L)^{2}\geq(-K_{X})^{2}L^{2}\geq 2.$
So $-K_{X}\cdot L\geq\sqrt{2}$. Also note that $-K_{X}\cdot L$ is a positive integer. Thus we have $-K_{X}.L\geq 2$. Therefore by [13, Theorem III (a)], $L$ is base point free. So $\varepsilon(X_{r},L,x)\geq 1$ for all $x\in X_{r}$. ∎
Corollary 3.4.
For $r\leq 7$, let $L$ be an ample line bundle on $X_{r}$. Then
$\varepsilon(X_{r},L,x)\geq 1$
for all $x\in X_{r}$.
Next, we consider the $r=8$ case. We first prove the following lemma.
Lemma 3.5.
Let $X_{8}$ be a blow up of $\mathbb{P}^{2}$ at 8 distinct points. If $L$ is an ample line bundle with $-K_{X_{8}}\cdot L=1$ then $L=-K_{X_{8}}$.
Proof.
Let $L=eH-n_{1}E_{1}-\dots-n_{8}E_{8}$ be an ample line bundle with $-K_{X_{8}}\cdot L=1$, i.e., $13e-\displaystyle\sum_{i=1}^{8}n_{i}=$ 1. If $L^{2}\geq 2$, the Hodge index theorem gives
$(L\cdot-K_{X_{8}})^{2}\geq L^{2}(-K_{X_{8}})^{2}\geq 2,$
which contradicts the hypothesis. So $L^{2}=1$, i.e., $e^{2}-\displaystyle\sum_{i=1}^{8}n_{i}^{2}=1$. So,
$89e^{2}-6e+1=(3e-1)^{2}=\bigg{(}\sum_{i=1}^{8}n_{i}\bigg{)}^{2}\leq 8\sum_{i=1}^{8}n_{i}^{2}=8e^{2}-$ 8 .
This reduces to $e^{2}-6e+9\leq 0$ which happens only if $e=3$. Since $L$ is ample we have $n_{i}>0$ and $3=e>n_{i}+n_{j}$ for all $i\neq j$. This gives $n_{i}=1$ for all $1\leq i\leq 8$. So $L=-K_{X_{8}}$. ∎
Theorem 3.6.
Let $X_{8}$ be a blow up of $\mathbb{P}^{2}$ at 8 distinct points $p_{1},p_{2},\dots,p_{8}$. Let $L$ be an ample line bundle on $X_{8}$. Suppose that $\varepsilon(X_{8},L,x)<1$ for some $x\in X_{8}$. Then the following must hold:
  1. (1)
    $p_{1},p_{2},\dots,p_{8}$ are smooth points on an irreducible singular cubic,
  2. (2)
    $L=-K_{X_{8}}$, and
  3. (3)
    $\varepsilon(X_{8},L,x)=\frac{1}{2}$.
Proof.
Suppose that $\varepsilon(X_{r},L,x)<1$ for an ample line bundle $L$ at some $x\in X_{r}$. By Proposition 3.1, the Seshadri constant is achieved by a curve $C$ on $X_{r}$ whose strict transform on $\widetilde{X}_{r}(x)$ is either a $(-1)$-curve or a $(-2)$-curve or a fixed component of $|-K_{\widetilde{X}_{r}(x)}|$ or a curve in the linear system $|-K_{\widetilde{X}_{r}(x)}|$. Let $C=dH-m_{1}E_{1}-\dots-m_{r}E_{r}$ and $t=\text{mult}_{x}C>1$, i.e,
$\varepsilon(X_{8},L,x)=\frac{L\cdot C}{\text{mult}_{x}C}=\frac{L\cdot C}{t}.$
First, we show that the strict transform of $C$ on $\widetilde{X}_{r}(x)$ is not a $(-1)$ or a $(-2)$-curve. If $C$ is such a curve, we have $C^{2}=t^{2}-i$, where $i=1,2$. Since
$\varepsilon(X_{8},L,x)=\frac{L\cdot C}{\text{mult}_{x}C}<1,$
we have $L\cdot C\leq t-1$. By the Hodge index theorem, $(L\cdot C)^{2}\geq L^{2}C^{2}\geq t^{2}-i$. Putting these together, we get
$(t-1)^{2}\geq(L\cdot C)^{2}\geq t^{2}-i.$
This is equivalent to $t\leq\frac{1+i}{2}$, where $i=1$ or $i=2$. This implies that $t=1$, which is a contradiction.
So $C$ is a singular curve whose strict transform is either a fixed component of $|-K_{\widetilde{X}_{r}(x)}|$ or a curve in the linear system $|-K_{\widetilde{X}_{r}(x)}|$. Since $C$ is a singular curve, we must have $d=3$. Hence $\widetilde{C}\in|-K_{\widetilde{X}_{r}(x)}|$. So $p_{1},\dots,p_{r}$ belong to the image of $C$ in $\mathbb{P}^{2}$. As $C$ is irreducible, it is clear that $\text{mult}_{x}C=2$. Since $\varepsilon(X_{8},L,x)<1$, the only possibility is
$\varepsilon(X_{8},L,x)=\frac{L\cdot C}{2}=\frac{1}{2}.$
Hence $L\cdot-K_{X_{8}}=1$ and by Lemma 3.5 this happens only if $L=-K_{X_{8}}$. ∎
Remark 3.7.
A part of Theorem 3.6 has an alternate proof using Theorem 2.1. Note that there exists a cubic passing through any distinct 8 points in $\mathbb{P}^{2}$. Theorem 2.1 says that $\frac{1}{2}$ is a Seshadri constant of $L$ only if $L\cdot-K_{X_{8}}=1$ and Lemma 3.5 enables us to conclude that it can happen only if $L=-K_{X_{8}}$.
Example 3.8.
By Theorem 3.6, the least Seshadri constant that can be achieved on $X_{8}$ is $\frac{1}{2}$. We now give an example in which $\frac{1}{2}$ is achieved at some point on a del Pezzo surface. Consider the smooth rational surface obtained by blowing up 8 smooth points $p_{1},p_{2},\dots p_{8}$ on a singular irreducible cubic in $\mathbb{P}^{2}$ such that no three lie on a line, no six lie on a conic, and if a singular cubic passes through all the points then $p_{i}$ are smooth points of that cubic for $1\leq i\leq 8$. Then $-K_{X_{8}}$ is ample by [4, Théorème 1]. Let $x$ denote the point in $X_{8}$ corresponding to the singular point on the cubic. Let $C$ denote the strict transform of the singular cubic containing the 8 points. From Theorem 3.6 it is clear that $\varepsilon(X_{8},-K_{X_{8}},x)\geq\frac{1}{2}$. Note that $\text{mult}_{x}C=2$ and $-K_{X_{8}}\cdot C=1$. These give us:
$\varepsilon(X_{8},-K_{X_{8}},x)=\frac{1}{2}.$
Proposition 3.9.
For $r\leq 5$, let $X_{r}$ denote the blow up of $r$ distinct points in $\mathbb{P}^{2}$, then $\varepsilon(X_{r},L,x)$ is an integer for all ample line bundles $L$ on $X_{r}$ and for all $x\in X_{r}$.
Proof.
It is known that the number of $(-1)$ and $(-2)$-curves in $\widetilde{X}_{r}(x)$ is finite for $r\leq 7$ and we can see that curves whose strict transforms are $(-1)$ or $(-2)$-curves are smooth in $\widetilde{X}_{r}(x)$ for $r\leq 5$. The curves in $X_{r}$ whose strict transform is a fixed component of $|-K_{\widetilde{X}_{r}(x)}|$ and is not an element of $|-K_{\widetilde{X}_{r}(x)}|$ are also smooth. It is also clear that a curve $C$ in $X_{r}$ such that $\widetilde{C}$ is an element of $|-K_{\widetilde{X}_{r}(x)}|$ has non-negative self intersection if $r\leq 5$. Note the fact that if $C$ is a Seshadri curve, its strict transform $\widetilde{C}$ has negative self intersection in $\widetilde{X}_{r}(x)$. This shows that all possible Seshadri curves are smooth when $r\leq 5$ which in fact implies that the Seshadri constant of every ample line bundle at every point of $X_{r}$ is an integer when $r\leq 5$. ∎
Remark 3.10.
It is shown in [2, Theorem 1.3] that when $r=6$, a fractional Seshadri constant is achieved. More precisely, $\varepsilon(X_{6},-K_{X_{6}},x)=\frac{3}{2}$ for some $x\in X_{6}$.
4. General blow ups of $\mathbb{P}^{2}$
We proceed with a similar setting as in the above section. Furthermore throughout this section we assume that $p_{1},p_{2},\dots,p_{r}$ are very general points in $\mathbb{P}^{2}$, where $r\geq 1$. We will give some lower bounds on Seshadri constants assuming the Strong SHGH Conjecture (Conjecture 1.3) is true.
Lemma 4.1.
For any point $x\in X_{r}$, let $\widetilde{X}_{r}(x)$ be the blow up of $X_{r}$ at x. Assume that Conjecture 1.3 is true. Let $C$ be a reduced, irreducible curve on $X_{r}$ and let $\widetilde{C}$ denote its strict transform on $\widetilde{X}_{r}(x)$. If $\widetilde{C}^{2}<0$ then $\widetilde{C}$ is one of the following types.
  1. (1)
    $\widetilde{C}^{2}=-1$ and $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}=-1$,
  2. (2)
    $\widetilde{C}^{2}=-1$ and $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}=1$, or
  3. (3)
    $\widetilde{C}^{2}=-2$ and $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}=0$
Proof.
Let $C=dH-m_{1}E_{1}-\dots-m_{r}E_{r}$ and $t=\text{mult}_{x}C$. So $\widetilde{C}=dH-m_{1}E_{1}-\dots-m_{r}E_{r}-tE_{x}$, where $E_{x}$ is the exceptional divisor in $\widetilde{X}_{r}(x)$. If $d=0$, $\widetilde{C}=E_{i}$ or $\widetilde{C}=E_{i}-E_{x}$. Both are of the required type.
Now let $d\geq 1$. Since $C$ is reduced, Conjecture 1.3 implies
(${d+2\choose 2}-\displaystyle\sum_{i=1}^{r}{m_{i}+1\choose 2}>{t+1\choose 2}-2.$
After some computations and rearrangements, the above inequality reduces to
$\frac{d^{2}-\displaystyle\sum_{i=1}^{r}m_{i}^{2}-t^{2}+3d-\sum_{i=1}^{r}m_{i}-t}{2}+1>-2,$
which is the same as
$\frac{\widetilde{C}^{2}-K_{\widetilde{X}_{r}(x)}.\widetilde{C}}{2}+1>-2.$
The Reimann-Roch theorem gives,
$h^{0}(\widetilde{X}_{r}(x),\widetilde{C})-h^{1}(\widetilde{X}_{r}(x),\widetilde{C})+h^{2}(\widetilde{X}_{r}(x),\widetilde{C})=\frac{\widetilde{C}^{2}-K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}}{2}+1.$
As $\widetilde{C}$ is an effective divisor in a smooth projective surface $\widetilde{X}_{r}(x)$, we have $h^{2}(\widetilde{X}_{r}(x),\widetilde{C})=0$. Since $\widetilde{C}^{2}<0$, $h^{0}(\widetilde{X}_{r}(x),\widetilde{C})=1$. Then
$-h^{1}(\widetilde{X}_{r}(x),\widetilde{C})=\frac{\widetilde{C}^{2}-K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}}{2}>-2.$
Hence $h^{1}(\widetilde{X}_{r}(x),\widetilde{C})\leq 1.$ We consider the following two cases:
  • Case 1:
    $h^{1}(\widetilde{X}_{r}(x),\widetilde{C})=0$.
    Then $h^{1}(\widetilde{X}_{r}(x),\widetilde{C})=\frac{K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}-\widetilde{C}^{2}}{2}=0$. So $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}=\widetilde{C}^{2}$.
    As $\widetilde{C}$ is reduced and irreducible, we can apply the adjunction formula:
    $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}+\widetilde{C}^{2}\geq-2.$
    Since $\widetilde{C}^{2}<0$, we get $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}=\widetilde{C}^{2}=-1.$
  • Case 2:
    $h^{1}(\widetilde{X}_{r}(x),\widetilde{C})=1$.
    Then $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}-\widetilde{C}^{2}=2$. As above, using the adjunction formula and $\widetilde{C}^{2}<0$, we conclude that $\widetilde{C}^{2}=-1$ and $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}=1$ or $\widetilde{C}^{2}=-2$ and $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}=0$.
The following is the main result of this section.
Theorem 4.2.
Assume that Conjecture 1.3 is true. Let $\pi:X_{r}\to\mathbb{P}^{2}$ be the blow up of $\mathbb{P}^{2}$ at very general points $p_{1},p_{2},\dots,p_{r}$. Then for every ample line bundle L on $X_{r}$ and for all $x\in X_{r}$,
$\varepsilon(X_{r},L,x)\geq 1.$
Proof.
Suppose that $\varepsilon(X_{r},L,x)<1$ for some ample line bundle $L$ and some $x\in X_{r}$. Then the Seshadri constant is achieved by an irreducible and reduced curve $C\subseteq X_{r}$, i.e.,
$\varepsilon(X_{r},L,x)=\frac{L\cdot C}{\text{mult}_{x}C}.$
Let $L=eH-n_{1}E_{1}-\dots-n_{r}E_{r}$ be an ample line bundle on $X_{r}$ and let $\pi_{x}:\widetilde{X}_{r}(x)\to X_{r}$ denote the blow up of $X_{r}$ at $x$. We know that $\widetilde{C}^{2}<0$, where $\widetilde{C}$ the is strict transform of $C$. Let $t=\text{mult}_{x}C$. Since $\varepsilon(X_{r},L,x)<1$ we have $t>1$ and $L\cdot C\leq t-1$. By Lemma 4.1, $20>\widetilde{C}^{2}\geq-$ 2 so $C^{2}=t^{2}-i$, where $i=1$ or $2$. Thus using the Hodge index theorem,
$(t-1)^{2}\geq(L\cdot C)^{2}\geq L^{2}C^{2}\geq t^{2}-i.$
So $(t-1)^{2}\geq t^{2}-i$, which reduces to $t\leq\frac{1+i}{2}$ for $i=1,2$. But this happens only if $t=1$, which is a contradiction. ∎
Remark 4.3.
From the above theorem, it follows that if Conjecture 1.3 is true then $\varepsilon(X_{r},x)$ and $\varepsilon(X_{r})$ cannot be zero for blows ups of a finite number of very general points in $\mathbb{P}^{2}$. A weaker assumption than Conjecture 1.3 gives the same conclusion, as we show below.
Conjecture 4.4.
Consider $X_{r}$ as in Theorem 4.2 and suppose that $d\geq 1$, $t\geq 2$ and $m_{1},m_{2},\dots,m_{r}\geq 0$ be integers such that
(${d+2\choose 2}-\displaystyle\sum_{i=1}^{r}{m_{i}+1\choose 2}\leq{t\choose 2}.$
Then any curve $C\in|dH-m_{1}E_{1}-\dots-m_{r}E_{r}|$ which has a point of multiplicity t is non-reduced.
Note that for $t\geq 2$, (${t+1\choose 2}-2\geq{t\choose 2}$. Hence Conjecture 4.4 follows from Conjecture 1.3.
Proposition 4.5.
Let $X_{r}$ be as in Theorem 4.2. If Conjecture 4.4 is true then $\varepsilon(X_{r},L,x)\geq\frac{1}{2}$ for all ample line bundles $L$ on $X_{r}$ and for any $x\in X_{r}$.
Proof.
Let $\pi_{x},\widetilde{X}_{r}(x),E_{x}$ be as above and let $L=eH-n_{1}E_{1}-n-\dots-n_{r}E_{r}$. To show that $\varepsilon(X_{r},L,x)\geq\frac{1}{2}$, it is enough to show that $\pi_{x}^{\ast}L-\frac{1}{2}E_{x}$ is nef, i.e., $2\pi_{x}^{\ast}L-E_{x}$ is nef. Equivalently, for any reduced, irreducible curve $\widetilde{C}=dH-m_{1}E_{1}-\dots-m_{r}E_{r}-tE_{x}$ in $\widetilde{X}_{r}(x)$, we need to show that
$2(de-m_{1}n_{1}-\dots-m_{r}n_{r})-t\geq 0$
If $t=1$ clearly $2(de-m_{1}n_{1}-\dots-m_{r}n_{r})-t\geq 0$ If $\widetilde{C}=E_{i}\text{ or }\widetilde{C}=E_{i}-E_{x}$ clearly $(\pi_{x}^{\ast}L-\frac{1}{2}E_{x})\cdot\widetilde{C}\geq 0$.
Now assume $d\geq 1$ and $t\geq 2$. Since $\widetilde{C}$ is a reduced curve in $\widetilde{X}_{r}(x)$, $C=\pi_{x}(\widetilde{C})$ is a reduced curve in $X_{r}$. So using Conjecture 4.4 we conclude:
(${d+2\choose 2}-\displaystyle\sum_{i=1}^{r}{m_{i}+1\choose 2}>{t\choose 2}.$
Consider a very general point $y\in X_{r}$ and let $p_{r+1}:=\pi(y).$ Therefore, by dimension count, there is a curve $D$ in $\mathbb{P}^{2}$ such that $\text{mult}_{p_{i}}D\geq m_{i}$ for all $i=1,2,\dots,r$ and $\text{mult}_{p_{r+1}}D\geq t-1$.
By [5, Theorem] we have, $\varepsilon(X_{r},L,y)\geq 1$. From this we conclude
$de-m_{1}n_{1}-\dots-m_{r}n_{r}-(t-1)\geq 0$
and ampleness of $L$ gives
$de-m_{1}n_{1}-\dots-m_{r}n_{r}\geq 1.$
Taking the sum of these two inequalities, we get
$2(de-m_{1}n_{1}-\dots-m_{r}n_{r})-t\geq 0$
This gives the required inequality
$\varepsilon(X_{r},L,x)\geq\frac{1}{2}.$
One of the most important questions in the area of Seshadri constants concerns their irrationality, i.e., is there an ample line bundle on a smooth surface so that at some point in the surface, the Seshadri constant of that line bundle is irrational? This will be addressed in a later section. But now we propose another question:
Question 4.6.
For a given $\frac{p}{q}\in\mathbb{Q}$, is there a positive integer $r$ and a line bundle $L$ on $X_{r}$ such that $\varepsilon(X_{r},L,x)=\frac{p}{q}$ for some $x\in X_{r}$.
Clearly Theorem 4.2 implies that if Conjecture 1.3 is true, the above question has a negative answer for $\frac{p}{q}<1$. So it is enough to look at $\frac{p}{q}$ with $p\geq q$. For this, it is enough to look for line bundles with Seshadri constant $\frac{p}{q}$, where $p$ is a prime number.
We give below a partial result in this direction.
Proposition 4.7.
For any $r$, there is an ample line bundle $L$ and $x\in X_{r}$ such that $\varepsilon(X_{r},L,x)=1$. Further, every positive integer is achieved as a Seshadri constant for some ample line bundle.
Proof.
By [7, Proposition 3.2], there exist a $d\in\mathbb{N}$ such that $L=dH-E_{1}-\dots-E_{r}$ is a very ample line bundle on $X_{r}$. Since $L$ is very ample we have $\varepsilon(X_{r},L,x)\geq 1$ for all $x\in X_{r}$. Note that $L\cdot E_{i}=1$ implies $\varepsilon(X_{r},L,x)\leq 1$ for all $x\in E_{i}$ for every $1\leq i\leq r$. So $\varepsilon(X_{r},L,x)=1$ for all $x\in\displaystyle\cup_{i=1}^{r}E_{i}$. ∎
5. Bounded Negativity Conjecture and Seshadri Constants
In the above sections, we looked at lower bounds for the Seshadri constants of ample line bundles on blow ups of $\mathbb{P}^{2}$. In this section, we will do the same for a surface more generally. As in Section 4 we assume some conjectures and come up with a lower bound for the Seshadri constant at any point on the surface.
Let $X$ be a surface and let $x\in X$. We assume the bounded negativity conjecture (Conjecture 1.5) on the blow up of $X$ at $x$ and prove a lower bound for Seshadri constants of ample line bundles on $X$ at $x$.
Theorem 5.1.
Let $X$ be a smooth complex projective surface. Let $x\in X$ and assume that Conjecture 1.5 is true for $\widetilde{X}(x)$, the blow up of $X$ at $x$, i.e., there exists a positive integer $b(\widetilde{X}(x))$ such that $C^{2}\geq-b(\widetilde{X}(x))$ for all reduced and irreducible curves $C$ in $\widetilde{X}(x)$. Then
$\varepsilon(X,x)\geq\frac{2}{b(\widetilde{X}(x))+1}.$
Proof.
Suppose that $\varepsilon(X,L,x)<\frac{2}{b(\widetilde{X}(x))+1}$ for an ample line bundle $L$ on $X$. Let $b_{x}=b(\widetilde{X}(x))$. Since $\frac{2}{b_{x}+1}\leq 1$ we have $\varepsilon(X,L,x)<1\leq\sqrt{L^{2}}$. So there is an irreducible and reduced curve $C\subset X$ such that
$\varepsilon(X,L,x)=\frac{L\cdot C}{\text{mult}_{x}C}.$
Let $t=\text{mult}_{x}C$. It is clear that $t>1$. Let $\widetilde{C}$ denote the strict transform of $C$ on $\widetilde{X}(x)$. By assumption, $\widetilde{C}^{2}=C^{2}-t^{2}\geq-b_{x}$. Since $L$ is an ample line bundle we have $L^{2}\geq 1$ and by the Hodge index theorem,
$(L\cdot C)^{2}\geq L^{2}C^{2}\geq t^{2}-b_{x}.$
Note that $\frac{L\cdot C}{t}=\varepsilon(X_{r},L,x)<1.$ So we have $L\cdot C\leq t-1$. Using these inequalities we conclude:
$(t-1)^{2}\geq(L\cdot C)^{2}\geq t^{2}-b_{x}.$
So
$(t-1)^{2}\geq t^{2}-b_{x},$
which is equivalent to
$\frac{1}{t}\geq\frac{2}{b_{x}+1}.$
This gives
$\varepsilon(X,L,x)=\frac{L\cdot C}{t}\geq\frac{1}{t}\geq\frac{2}{b_{x}+1},$
which is a contradiction. ∎
In earlier sections, we also found a lower bound for $\varepsilon(X)$, when $X$ is a blow up of $\mathbb{P}^{2}$. In general, we need a stronger assumption on $X$ to conclude that $\varepsilon(X)>0$. For that, we formulate a stronger conjecture as follows:
Conjecture 5.2.
Let $X$ be a smooth complex projective surface. Then there exists an integer $B(X)>0$ such that if $C$ is an irreducible, reduced curve on $X$, the strict transform of $C$ on any 1- point blow up of $X$ has self-intersection at least $-B(X)$, i.e., if $\widetilde{C}$ denotes the strict transform of $C$ then
$\widetilde{C}^{2}\geq-B(X).$
Theorem 5.3.
Assume that Conjecture 5.2 is true for a smooth projective surface $X$. Then
$\varepsilon(X)\geq\frac{2}{B(X)+1}.$
Proof.
Note that $b(\widetilde{X}(x))\leq B(X)$. So, using the proof of Theorem 5.1, we can show that
$\varepsilon(X,x)\geq\frac{2}{B(X)+1}$
for all $x\in X$. Hence
$\varepsilon(X)\geq\frac{2}{B(X)+1}.$
Now we will give certain surfaces for which the Conjecture 5.2 is true.
Example 5.4.
Throughout this paper, we are interested in $\mathbb{P}^{2}$ and blow ups of $\mathbb{P}^{2}$ at finitely many points in $\mathbb{P}^{2}$. It is easy to see that every negative self-intersection curve in a 1-point blow up of $\mathbb{P}^{2}$ is $(-1)$-curve. So Conjecture 5.2 is true for $B(\mathbb{P}^{2})=1$. Let $X_{r}$ be the blow up of $r$ very general points in $\mathbb{P}^{2}$. Then Lemma 4.1 implies that if Conjecture 1.3 is true then Conjecture 5.2 is true for $B(X_{r})=-2$.
Example 5.5.
Consider the surface $X_{r}$ which is a blow up of $\mathbb{P}^{2}$ at $r\leq 8$ points such that no three points lie on a line, no six points lie on a conic and no eight points lie on a singular cubic. So by Proposition 3.3 and Theorem 3.6, we know that the anticanonical divisor of every one point blow up of $X_{r}$ is nef, i.e., for every $x\in X_{r}$ the anticanonical divisor $-K_{\widetilde{X}_{r}(x)}$ is a nef divisor. So, $-K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}\geq 0$ for every irreducible reduced curve $C$ in $X_{r}$. Now using the adjunction formula we have
$-2\leq 2P_{a}(\widetilde{C})-2=\widetilde{C}^{2}+K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}\leq\widetilde{C}^{2}.$
This guarantees that for $r\leq 8$, Conjecture 5.2is true with $B(X_{r})=-2$
6. Submaximality of Line Bundles on General Blow Ups of $\mathbb{P}^{2}$
Let $X_{r}$ denote a blow up of $r$ distinct points in $\mathbb{P}^{2}$. Let $\text{Nef}(X_{r})$ denote the nef cone of $X_{r}$. Fix an $x\in X_{r}$. In this section, we are interested in the Seshadri function
$\varepsilon_{x}:\text{Nef}(X_{r})\rightarrow\mathbb{R}$
defined as follows:
$\varepsilon_{x}(L)=\varepsilon(X_{r},L,x).$
It is known that $\varepsilon_{x}$ is continuous; see [18]. For an abelian surface, [18, Lemma 4.1] shows that $\varepsilon_{x}$ is piece-wise linear except for a nowhere dense set. We study this function for $X_{r}$ in this section.
Lemma 6.1.
Let $X_{r}$ be a blow up of r distinct points in $\mathbb{P}^{2}$ and let $L=eH-n_{1}E_{1}-\dots-n_{r}E_{r}$ be an ample line bundle on $X_{r}$. If $e>\sum_{i=1}^{r}n_{i}^{2}$ then $\varepsilon(X_{r},L,x)<\sqrt{L^{2}}$, for all $x\in X_{r}$.
Proof.
Clearly $H-E_{1}$ is an effective divisor on $X_{r}$. So
$e-n_{1}=L\cdot(H-E_{1})>0.$
So $en_{1}>n_{1}^{2}$. Note that $n_{1}=L\cdot E_{1}>0$. Now if $e>\sum_{i=1}^{r}n_{i}^{2}$, we have $en_{1}\geq e>\sum_{i=1}^{r}n_{i}^{2}$. Hence $2n_{1}e>\big{(}\sum_{i=1}^{r}n_{i}^{2}\big{)}+n_{1}^{2}$ This implies $L\cdot(H-E_{1})<\sqrt{L^{2}}$. So $\varepsilon(X_{r},L,x)<\sqrt{L^{2}}$. ∎
Proposition 6.2.
Let $X_{r}$ be a blow up of r distinct points in $\mathbb{P}^{2}$. For each $x\in X_{r}$, there exists a non-empty open set $U\subset\rm{Nef}(X_{r})$ such that $\varepsilon(X_{r},L,x)<\sqrt{L^{2}}$ for all $L\in U$.
Proof.
Let
$S=\{L\in\text{Nef}(X_{r})~{}~{}|~{}~{}L=eH-n_{1}E_{1}-\dots-n_{r}E_{r}\text{ with }e>\sum_{i=1}^{r}n_{i}^{2}\}.$
By Lemma 6.1, $U=S\cap\text{Amp}(X_{r})$ is an open set in $\text{Nef}(X_{r})$ such that $\varepsilon(X_{r},L,x)<\sqrt{L^{2}}$ for all $L\in U$. It remains to prove that $U$ is non-empty. $L=dH-n_{1}E_{1}-\dots-n_{r}E_{r}$ be an ample line bundle. It is clear that $eH-n_{1}E_{1}-\dots-n_{r}E_{r}$ is ample $\forall e>d$. Hence $S\cap\text{Amp}(X_{r})$ is non-empty. ∎
For an abelian surface $X$ and $x\in X$, the set
$W_{x}=\{L\in\text{Nef}(X)~{}~{}|~{}~{}\varepsilon(L,x)=\sqrt{L^{2}}\}$
is nowhere dense [18, Lemma 4.1]. We can ask a similar question for $X_{r}$. Let $W_{x}$ denote the same set for $X_{r}$. The following propositions show conjecturally that $W_{x}$ is not nowheredense in $X_{r}$.
We say that a line bundle $F=eH-n_{1}E_{1}-\dots-n_{r}E_{r}$ is in standard form if $n_{1}\geq n_{2}\geq\dots\geq n_{r}\geq 0$ and $d\geq n_{1}+n_{2}+n_{3}$. For a line bundle $F$ in standard form, it is well-known that $F\cdot C\geq 0$ for all $(-1)$-curves $C$. Following lemma is a generalization of this fact.
Lemma 6.3.
Let $L=eH-n_{1}E_{1}-\dots-n_{r}E_{r}$ be a line bundle on $X_{r}$ with $n_{i}\geq 0$ and $e\geq n_{i}+n_{j}+n_{k}$ for all distinct $i,j,k\in\{1,2,\dots,r\}$. Then $L\cdot C\geq 0$ for every $(-1)$-curve $C$
Proof.
If $C$ is any of the exceptional divisors $E_{1},\dots,E_{r}$ then $L\cdot C=L\cdot E_{i}=n_{i}\geq 0$. Now assume that $C$ is a $(-1)$-curve and $C\neq E_{i}$ for $1\leq i\leq r$. Let $S_{r}$ denote the symmetric group on $\{1,2,\dots,r\}$. Choose $\sigma\in S_{r}$ such that $n_{i}^{\prime}\geq n_{i+1}^{\prime}$, where $n_{i}^{\prime}=n_{\sigma(i)}$. Let $L^{\prime}=eH-n_{1}^{\prime}E_{1}^{\prime}-\dots-n_{r}^{\prime}E_{r}^{\prime}$ where $E_{i}^{\prime}=E_{\sigma(i)}$. Then we can write $L^{\prime}$ as a non-negative integer combination of
$H,H_{1}=H-E_{1}^{\prime},H_{2}=2H-E_{1}^{\prime}-E_{2}^{\prime},H_{i}=3H-E_{1}^{\prime}-E_{2}^{\prime}-\dots-E_{i}^{\prime},~{}~{}\forall 3\leq i\leq r$
as follows:
$L^{\prime}=(e-(n_{1}^{\prime}+n_{2}^{\prime}+n_{3}^{\prime}))H+(n_{1}^{\prime}-n_{2}^{\prime})H_{1}+(n_{2}^{\prime}-n_{3}^{\prime})H_{2}+(n_{3}^{\prime}-n_{4}^{\prime})H_{3}+\dots+(n_{r-1}^{\prime}-n_{r}^{\prime})H_{r-1}+n_{r}^{\prime}H_{r}.$
Note that for $3\leq i\leq r$, $H_{i}+K_{X_{r}}=E_{i+1}^{\prime}+\dots+E_{r}^{\prime}$. As $C\neq E_{i}$, we have $C\cdot E_{i}^{\prime}\geq 0$. So $H_{i}\cdot C\geq-K_{X_{r}}\cdot C$. Note that $H,H_{1},H_{2}$ are nef and since $C$ is a $(-1)$-curve, we have $-K\cdot C=1$. So $L^{\prime}\cdot C\geq 0$. Now we are done since $L\cdot C=L^{\prime}\cdot C$. ∎
From now onwards $X_{r}$ will denote a blow up of $r$ very general points in $\mathbb{P}^{2}$. We will now prove some result about the Seshadri function for a very general point $x\in X_{r}$ assuming Weak SHGH Conjecture (Conjecture 1.4).
Theorem 6.4.
For $r\geq 9$, assume that Conjecture 1.4 is true. Then for a very general point $x\in X_{r}$, there exists an open set of $\rm{Nef}(X_{r})$ on which $\varepsilon_{x}$ is not linear. Moreover, $\varepsilon_{x}$ is not a piecewise linear function.
Proof.
Since $L\mapsto\sqrt{L^{2}}$ is not a linear function on $\text{Nef}(X_{r})$, it is enough to construct an open set $U_{x}$ of $\text{Nef}(X_{r})$ such that $\varepsilon(X_{r},L,x)=\sqrt{L^{2}}$ for all line bundles $L\in U_{x}$.
Consider the following subset of $\text{NS}_{\mathbb{R}}(X_{r})$,
$V_{x}=\{L=eH-n_{1}E_{1}-\dots-n_{r}E_{r}~{}~{}|~{}~{}\sqrt{L^{2}}>n_{j}\text{ and }e>\sqrt{L^{2}}+n_{j}+n_{k}~{}~{}\forall 1\leq j<k\leq r\}.$
Clearly $V_{x}$ is open in $\text{NS}_{\mathbb{R}}(X_{r})$. Now we will prove that $U_{x}=V_{x}\cap\text{Amp}(X_{r})$ has the desired property. Consider $L=eH-m_{1}E_{1}-\dots-m_{r}E_{r}\in U_{x}$ with $e,m_{1},\dots,m_{r}\in\mathbb{Q}$. If $\varepsilon(L,x)<\sqrt{L^{2}}$, then $\varepsilon(nL,x)<\sqrt{(nL)^{2}}$ for any positive integer $n$. Choose $n\in\mathbb{N}$ such that $nL$ is an integral line bundle. So there is a curve $C$ containing $x$ such that $\frac{nL\cdot C}{\text{mult}_{x}C}=\varepsilon(nL,x)$. The same $C$ will serve as a Seshadri curve for $L$ also.
Now we argue as in the proof of [10, Proposition 2.1]. Let $\pi_{x}:\widetilde{X}_{r}(x)\rightarrow X_{r}$ be the blow up of $X_{r}$ at $x$. As $C$ is a Seshadri curve, $\widetilde{C}^{2}<0$, where $\widetilde{C}$ is strict transform of $C$. Since we assumed that Conjecture 1.4 is true, $\widetilde{C}$ is a $(-1)$-curve. Then
$0=(\pi_{x}^{\ast}(L)-\varepsilon(L,x)E)\cdot\widetilde{C}>(\pi_{x}^{\ast}(L)-(\sqrt{L^{2})}E)\cdot\widetilde{C}.$ (6.1)
By the choice of $L$, we have $\sqrt{L^{2}}>n_{i}$ and $e>\sqrt{L^{2}}+n_{i}+n_{j}$, $\forall 1\leq i\neq j\leq r$. Therefore by Lemma 6.3, we have $(\pi_{x}^{\ast}(L)-\sqrt{L^{2}}E)\cdot\widetilde{C}\geq 0$ for all $(-1)$-curves $C$. This is a contradiction to (6.1). Hence $\varepsilon(L,x)=\sqrt{L^{2}}$ for every $L\in U_{x}\cap\text{NS}_{\mathbb{Q}}(X_{r})$. So the continuous function $\varepsilon_{x}(L)$ is equal to $\sqrt{L^{2}}$ for all $L$ in the dense set $U_{x}\cap\text{NS}_{\mathbb{Q}}(X_{r})$. Hence the Seshadri constant is maximum for all line bundles in $U_{x}$.
It remains to prove that $U_{x}$ is non-empty. For $r>12$ and $r\neq 15,16$, it is easy to show that there exists $e$ such that $14e-4<r<e^{2}-$ 1. Consider $L=eH-E_{1}-\dots-E_{r}$ for such an $e$. It is easy to see that the Nagata conjecture implies that a uniform line bundle $L$ is ample if $L^{2}>0$. Note that Conjecture 1.4 implies the Nagata Conjecture. Since $\sqrt{L^{2}}=\sqrt{e^{2}-r}>0$, $L$ is an ample line bundle. Now to say $L\in U_{x}$ is enough to prove $e>\sqrt{e^{2}-r}+2$, which is equivalent to $r>4e-4$. So for $r>12$,   $r\neq 15,16$ the theorem is true. For $r\in\{9,10,11,12,15,16\}$, we explicitly give an element in $U_{x}$ below:
$r=9~{}$ : ${}16H-5E_{1}-\dots-5E_{r}$,
$r=10$ : $10H-3E_{1}-\dots-3E_{r}$,
$r=11$ : $7H-2E_{1}-\dots-2E_{r}$,
$r=12$ : $11H-3E_{1}-\dots-3E_{r}$,
$r=15$ : $9H-2E_{1}-\dots-2E_{r}$,
$r=16$ : $9H-2E_{1}-\dots-2E_{r}$,
We can easily check that the above line bundles satisfy the conditions to be in $U_{x}$. For example, consider the case $r=9$. Let $L=16H-5E_{1}-\dots-5E_{r}$, then $L$ is in $U_{x}$ as $\sqrt{31}=\sqrt{L^{2}}>n_{i}=5$ and $516=e>\sqrt{L^{2}}+n_{j}+n_{k}=\sqrt{31}+5+$ 5. Other cases are similar. ∎
Now consider $W_{x}^{\prime}=\{L\in\text{Nef}(X_{r})~{}~{}|~{}~{}\varepsilon(X_{r},L,x)\notin\mathbb{Q}\}$. We ask if we can have infinitely many line bundles in $W_{x}^{\prime}$, which are not integer multiples of each other?
Corollary 6.5.
Assume that Conjecture 1.4 is true. For $r\geq 9$, there are infinitely many integral ample line bundles, which are not integer multiples of each other on $X_{r}$ with irrational Seshadri constant at a very general point $x\in X_{r}$.
Proof.
In the proof of Theorem 6.4, for $r\geq 13$ and $r\neq 15,16$, we have constructed an open set of $\text{Nef}(X_{r})$ inside $W_{x}$ containing a uniform line bundle $dH-E_{1}-\dots-E_{r}$ for some $d\in\mathbb{N}$. So $(d+\frac{1}{n})H-E_{1}-\dots-E_{r}$ is in $W_{x}$ for large $n$. We have $\varepsilon(X,nL,x)=n\varepsilon(X,L,x)$. Then $L_{n}=(dn+1)H-nE_{1}-\dots-nE_{r}$ has Seshadri constant equal to $\sqrt{L_{n}^{2}}$. So it is enough to prove that $\sqrt{L_{n}^{2}}$ is irrational for infinitely many $n$.
For $r\geq 13$ and $r\neq 15,16$, we can choose $d$ such that $104d-3\leq r\leq 6d-$ 10. Then we conclude that
$(dn-3n)^{2}+(n^{2}+2dn+1)\leq(dn+1)^{2}-rn^{2}=L_{n}^{2}\leq(dn-2n)^{2}-(n^{2}+2dn+1).$
This implies that there are at most $n$ possibilities for $L_{n}^{2}$ to be a perfect square. To be more precise $L_{n}^{2}$ can be a prefect square only if for some $1\leq i\leq n$,
$(dn+1)^{2}-rn^{2}=(dn-3n+i)^{2}.$
This will reduce to a quadratic equation in $n$ as
$n^{2}(-r-9+6d)+n(2d-2id+6i)+1-i^{2}=0.$
So for a fixed $d,r,i$ there are only two choices for $n$, and there are finitely many options for $n$ also. So $L_{n}^{2}$ is not a perfect square except for finitely many $n$. It is easy to see that $L_{n}=L_{m}$ only if $n=m$. A similar calculation will give the result for $r=9,10,11,12,14,15,16$. ∎
Note that Corollary 6.5 strengthens [10, Theorem 2.4] since we exhibit infinitely many ample line bundles which have irrational Seshadri constants.
Now we exhibit some result regarding the Seshadri function for an arbitrary $x\in X_{r}$ assuming Strong SHGH Conjecture (Conjecture 1.3).
Lemma 6.6.
Assume that Conjecture 1.3 is true. For $x\in X_{r}$, let $\widetilde{X}_{r}(x)$ be the blow up of $X_{r}$ at $x$ and let $E_{x}$ denote the corresponding exceptional divisor. Let $L=eH-n_{1}E_{1}-\dots-n_{r}E_{r}-n_{x}E_{x}\in\rm{NS}_{\mathbb{R}}(\widetilde{X}_{r}(x))$ be a line bundle on $\widetilde{X}_{r}(x)$ with $e\geq n_{i}+n_{j}+2n_{k}$, $n_{i}\geq n_{x}$ and $n_{x}\geq 0$ for all distinct $i,j,k\in\{1,\dots,r\}$. Then $L\cdot C\geq 0$ for all irreducible, reduced curves $C$ with $C^{2}<0$.
Proof.
If $C=E_{i}$ for some $i\in\{1,2,\dots,r\}$ then $L\cdot C=L\cdot E_{i}=n_{i}\geq 0$. If $C=E_{i}-E_{x}$ for some $i\in\{1,2,\dots,r\}$ then $L\cdot C=L\cdot(E_{i}-E_{x})=n_{i}-n_{x}\geq 0$. Now assume that $C$ is any curve other than $E_{i}$ or $E_{i}-E_{x}$, with $C^{2}<0$.
Choose a permutation $\sigma\in S_{r}$ such that $n_{i}^{\prime}\geq n_{i+1}^{\prime}$ where $n_{i}^{\prime}=n_{\sigma(i)}$. Let $L^{\prime}=eH-n_{1}^{\prime}E_{1}^{\prime}-\dots-n_{r}^{\prime}E_{r}^{\prime}-n_{x}E_{x}$ where $E_{i}^{\prime}=E_{\sigma(i)}$. Now we can write $L^{\prime}$ as a non-negative integer combination of $H,H_{1}=H-E_{1}^{\prime},H_{1}=2H-E_{1}^{\prime}-E_{2}^{\prime},H_{i}=3H-E_{1}^{\prime}-\dots-E_{i}^{\prime}$, $\forall 3\leq i\leq r$ and $H_{x}=-K_{\widetilde{X}_{r}(x)}$ as below:
$L^{\prime}=(e-(n_{1}^{\prime}+n_{2}^{\prime}+n_{3}^{\prime}))H+(n_{1}^{\prime}-n_{2}^{\prime})H_{1}+(n_{2}^{\prime}-n_{3}^{\prime})H_{2}+\dots+(n_{r-1}^{\prime}-n_{r}^{\prime})H_{r-1}+(n_{r}^{\prime}-n_{x})H_{r}+n_{x}H_{x}.$
As we explained in proof of Lemma 6.3, $H_{i}\cdot C\geq-K\cdot C$ for $3\leq i\leq r$ and $H,H_{0},H_{1}$ are nef. Also note that $H\cdot C\geq 1$. By Lemma 4.1, $-K_{\widetilde{X}_{r}(x)}\cdot C\geq-1$. So we have $H_{i}\cdot C\geq-1$ for $3\leq i\leq r$. Therefore,
$\begin{split}L\cdot C&=L^{\prime}\cdot C\\ &=(e-(n_{1}^{\prime}+n_{2}^{\prime}+n_{3}^{\prime}))H\cdot C+(n_{1}^{\prime}-n_{2}^{\prime})H_{1}\cdot C+(n_{2}^{\prime}-n_{3}^{\prime})H_{2}\cdot C+\dots+\\ &\hskip 14.22636pt(n_{r}^{\prime}-n_{x})H_{r}\cdot C+n_{x}H_{x}\cdot C\\ &\geq(e-(n_{1}^{\prime}+n_{2}^{\prime}+n_{3}^{\prime}))+(n_{3}^{\prime}-n_{4}^{\prime})(-1)+\dots+(n_{r}^{\prime}-n_{x})(-1)+n_{x}(-1)\\ &\geq(e-(n_{1}^{\prime}+n_{2}^{\prime}+n_{3}^{\prime}))+(-n_{3}^{\prime})\\ &\geq 0.\end{split}$
The last inequality follows from the hypotheses. ∎
Theorem 6.7.
For $r\geq 16$, assume that Conjecture 1.3 is true. For any $x\in X_{r}$, there is a non-empty open set $U\subset\rm{Nef}(X_{r})$ such that $\varepsilon(X_{r},L,x)=\sqrt{L^{2}}$ for all $L\in U$.
Proof.
Consider the following subset of $\text{NS}_{\mathbb{R}}(X_{r})$,
$V_{x}=\{L=eH-n_{1}E_{1}-\dots-n_{r}E_{r}~{}~{}|~{}~{}\sqrt{L^{2}}<n_{i},e>n_{j}+n_{k}+2n_{l}\text{ for distinct }j,k,l\in\{1,\dots,r\}\}.$
Clearly $V_{x}$ is open in $\text{NS}_{\mathbb{R}}(X_{r})$. We will prove that $U_{x}=V_{x}\cap\text{Amp}(X_{r})$ has the desired property.
First, we will show $U_{x}$ is non-empty. For $r\geq 16$, consider $e\in\mathbb{Q}_{+}$ such that
$r(r-1)^{2}<e^{2}<(r+1)(r-1)^{2}\text{ and }L=eH-(r-1)E_{1}-\dots-(r-1)E_{r}.$
As before, we can show that $L$ is ample, since $L$ is uniform and $L^{2}>0$ and we have $e>\sqrt{r}(r-1)\geq 4(r-1)$. To show that $L\in U_{x}$ it remains to prove that $\sqrt{e^{2}-r(r-1)^{2}}<r-1$, which is equivalent to $e^{2}<(r+1)(r-1)^{2}$. This is true by the choice of $e$. So $U_{x}$ is non-empty.
Consider $L=eH-m_{1}E_{1}-\dots-m_{r}E_{r}\in U_{x}$ with $e,m_{1},\dots,m_{r}\in\mathbb{Q}_{+}$. Suppose that $\varepsilon(L,x)<\sqrt{L^{2}}$. The same is true for any integral multiple of $L$. Let $n$ be a positive integer such that $nL$ has integer coefficients. So there is a curve $C$ such that $\frac{nL\cdot C}{\text{mult}_{x}C}=\varepsilon(nL,x)$. The same $C$ will serve as a Seshadri curve for $L$ also.
Let $\pi_{x}:\widetilde{X}_{r}(x)\rightarrow X_{r}$ be the blow up of $X_{r}$ at $x$ and let $E_{x}$ denote the exceptional divisor. As $C$ is a Seshadri curve, we have $\widetilde{C}^{2}<0$, where $\widetilde{C}$ is the strict transform of $C$. Then
$0=(\pi_{x}^{\ast}(L)-\varepsilon(L,x)E_{x})\cdot\widetilde{C}>(\pi_{x}^{\ast}(L)-\sqrt{L^{2}}E_{x})\cdot\widetilde{C}.$ (6.2)
By the choice of $L$, $\sqrt{L^{2}}<n_{i}$ and $e>n_{i}+n_{j}+2n_{k}$ for distinct $i,j,k\in\{1,\dots,r\}$. Therefore by Lemma 6.6, we have $(\pi_{x}^{\ast}(L)-\sqrt{L^{2}}E)\cdot\widetilde{C}\geq 0$. This is a contradiction to (6.2). Hence for every $L\in U_{x}\cap\text{NS}_{\mathbb{Q}}(X_{r})$, $\varepsilon(L,x)=\sqrt{L^{2}}$. Since $U_{x}\cap\text{NS}_{\mathbb{Q}}(X_{r})$ is dense in $U_{x}$, the Seshadri constant is maximum for all line bundles in $U_{x}$. ∎
Acknowledgments
It is a pleasure to thank my advisor Krishna Hanumanthu for a meticulous reading of the manuscript and for many helpful discussions, suggestions, and comments.